# Four-dimensional phase space tomography from one-dimensional measurements in a high-power hadron ring

### v1

@@ Abstract
In this paper, we use one-dimensional measurements to infer the four-dimensional phase space density of an accumulated 1 GeV proton beam in the Spallation Neutron Source (SNS) accelerator. The reconstruction was performed using MENT, an exact maximum-entropy tomography algorithm, and thus represents the most reasonable inference from the data. The reconstructed distribution reproduces the measured profiles with the same dynamic range as the measurement devices, and simulations indicate that the problem is well-constrained. Similar measurements could serve as benchmarks for simulations of intense, coupled beam dynamics in the SNS or other hadron rings.
@@ I Introduction
The evolution of an intense hadron beam in an accelerator depends strongly on its initial distribution in six-dimensional phase space. There is typically uncertainty attached to the distribution because conventional diagnostics measure one or two dimensions at a time. Direct high-dimensional phase space imaging could address this problem in hadron linacs [1]. In hadron rings, however, the beam energy and intensity preclude direct measurements. High-dimensional information could nonetheless benefit certain applications in hadron rings. For example, proposed methods such as eigenpainting [2] and nonlinear integrable optics [3] generate intense beams with strong interplane correlations in coupled focusing systems. Four-dimensional information could also be used to reliably predict the two-dimensional density on high-power targets in operational accelerators.
An alternative to direct phase space imaging is phase space tomography, where the phase space density is inferred from low-dimensional views. Recent experiments have demonstrated four-dimensional tomography from two-dimensional measurements in electron accelerators [4, 5, 6], as well as in H- linacs using laser-wire diagnostics [7]. Two-dimensional measurements are not typically available in hadron rings, where the primary diagnostics are wire scanners, which record the beam density on a one-dimensional axis in the transverse plane. [^1] The two-dimensional beam image on a high-power target can be viewed by coating the target with a luminescent material. A long optical fiber system is required to view the images because of the high radiation levels near the target [8]. We did not attempt to infer the four-dimensional density from such images. Strict requirements on the beam shape on the target make it difficult to perform the necessary quadrupole scans. Additionally, the images are noisy and degrade over the target lifetime. It is not a priori obvious whether a realistic number of one-dimensional views can constrain a four-dimensional distribution, nor whether the required views can be obtained in a real machine.
In this paper, we use wire-scanner measurements to reconstruct the four-dimensional phase space density of an accumulated 1 GeV proton beam in the Spallation Neutron Source (SNS) accelerator. We follow the work of Minerbo, Sander, and Jameson [9], who applied four-dimensional tomography to one-dimensional data in a low-energy hadron linac nearly 45 years ago. The reconstruction utilizes cross-plane information from diagonal wires mounted alongside horizontal and vertical wires on each device. We reconstruct the distribution in two steps: first, we fit the (overdetermined) covariance matrix of second-order moments to the measured rms beam sizes; second, we incorporate the covariance matrix in a prior distribution and maximize the distribution’s entropy relative to this prior, subject to the measurement constraints. The reconstructed distribution reproduces the measured profiles with the same dynamic range as the measurement devices, and simulations indicate that the problem is well-constrained. Similar measurements could serve as benchmarks for simulations of intense, coupled beam dynamics in the SNS or other hadron rings.
II Experiment
Our experiment was conducted at the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory (ORNL), shown in Fig. 1. The SNS bombards a liquid mercury target with intense proton pulses at 60 Hz repetition rate [10]. Each pulse is the superposition of 1000 “minipulses” injected from a 400-meter linac into a 250-meter storage ring at 1.3 GeV kinetic energy. The final pulse contains approximately $1.4\times 10^{14}$ protons, corresponding to 1.7 MW beam power. The accumulated beam is extracted and transported from the ring to the target in a 150-meter Ring-Target Beam Transport (RTBT).
Figure 1: Layout of the Spallation Neutron Source (SNS) accelerator. H- ion minipulses are repeatedly accelerated to 1 GeV in the linac and converted to protons by passing through a carbon foil at the ring injection point. A 700 nanosecond proton pulse is accumulated over 1000 turns in the ring. Dipole magnets vary the mean phase space coordinates ($x$, $x^{\prime}$, $y$, $y^{\prime}$) of the circulating beam relative to the injected beam as a function of time during accumulation. In the Ring-Target Beam Transport (RTBT), four wire scanners measure the particle density on a horizontal, vertical, and diagonal axis ($x$, $y$, $u$).
Each wire scanner consists of three 100-$\mu$m tungsten wires mounted on a fork at forty-five-degree angles relative to each other. The secondary electron emission from each wire gives the beam density on the horizontal ($x$), vertical ($y$), and diagonal ($u$) axis. Each wire scanner operates at 1 Hz repetition rate; thus, each point in the measured profile corresponds to a different beam pulse. A single measurement takes approximately five minutes, including the return to the initial wire position; however, the four wire scanners run in parallel, so each scan generates twelve profiles.
Fig. 2 shows the locations of the wire scanners and nearby quadrupole magnets Fig. 2. Two power supplies control the first eight quadrupoles, while the remaining quadrupoles before the target have independent power supplies. We aim to reconstruct the four-dimensional phase space distribution at a point before the first varied quadrupole (QH18).
Given a finite number of measurements, it is not entirely clear how to find the optimal set of views—those that place the tightest constraints on the unknown four-dimensional density. Nor is it entirely clear how many views are required to reconstruct the phase space density with sufficient accuracy. These questions deserve further study, especially when the focusing system has coupled or nonlinear elements. Here, we draw on previous experiments in electron accelerators that attempted to reconstruct the four-dimensional phase space distribution from two-dimensional projections onto the $x$-$y$ plane in an uncoupled focusing system. Our focusing system is also uncoupled, and each wire scanner provides three one-dimensional projections that constrain the density in the $x$-$y$ plane. Thus, we hypothesize that observations in these experiments are relevant to us.
Hock and Wolski [11] discovered a near closed-form solution for the four-dimensional phase space density from a set of two-dimensional measurements. The solution assumes independent rotations in the $x$-$x^{\prime}$ and $y$-$y^{\prime}$ planes by the phase advances $\mu_{x}$ and $\mu_{y}$. In other words, in a normalized frame, the transfer matrix takes the form
Sampling the entire $\mu_{x}$-$\mu_{y}$ plane leads to an exact solution for $f(x,x^{\prime},y,y^{\prime})$. A lengthy nested scan utilizing this principle has been demonstrated [12]; however, accurate reconstructions appear to be possible without covering the entire range of phase advances. Examples include holding one phase advance fixed while the other varies [4] or varying both phases in a single quadrupole scan [6]. These studies suggest that $\mu_{x}$, $\mu_{y}$, and $\mu_{x}-\mu_{y}$ should cover a range of values between 00 and $\pi$.
The constraints in the SNS limit our control of the phase advances at each wire scanner. In [13], we found that the nominal optics produce little variation in $\mu_{x}-\mu_{y}$, leading to an ill-conditioned system when fitting the covariance matrix. We found a second set of optics that led to a better-conditioned problem; in the second set of optics, $\mu_{x}$ and $\mu_{y}$ at the last wire scanner (WS24) were shifted by 45 degrees in opposite directions relative to their nominal values. In the present study, we added a third set of optics with different phase advances at WS24. The phase advances at each wire scanner over the three sets of optics are plotted in Fig. 3. We found that this set of optics leads to accurate four-dimensional phase space density reconstructions when tested with known initial distributions.
We measured a beam generated by a nonstandard injection procedure intended to generate strong cross-plane correlations in the beam. (The standard SNS painting method naturally washes out cross-plane correlations during injection.) We collected beam profiles for three sets of optics in approximately fifteen minutes. Each of the 36 profiles was centered and thresholded to remove background noise. We modeled the accelerator lattice (which contains only drifts and quadrupoles) as a linear transformation of the phase space coordinates. We computed the transfer matrices from the OpenXAL [14] online model of the SNS accelerator.
III Reconstruction Results
As detailed in Appendix A, we performed the reconstruction in two steps. First, we fit the overdetermined covariance matrix to the measured rms beam sizes using Linear Least Squares (LLSQ). Second, we used this covariance matrix to define a Gaussian prior distribution and used the MENT algorithm [15] to maximize the entropy relative to this prior. The result is the most conservative inference from the data.
Fig. 4 shows the LLSQ fit of the covariance matrix to the measured rms beam size on each wire. The fit reproduces the rms beam sizes on all wires in all scan steps. The tightly fit ellipses over a range of phase advances in the $x$-$y^{\prime}$ and $y$-$y^{\prime}$ planes indicate a well-modeled lattice and small measurement errors. Standard LLSQ error propagation gives uncertainties of a few percent for the cross-plane covariances.
Figure 4: Least-squares covariance matrix fit. Top: vertical lines in $x$-$x^{\prime}$ and $y$-$y^{\prime}$ phase space mapped back to the reconstruction point and plotted in normalized coordinates. In these units, the best-fit ellipse is the unit circle. Bottom: measured (black) vs. fit (red) rms beam sizes on each wire, assuming $5\%$ measurement error. Bottom right: best-fit correlation matrix.
With a Gaussian prior based on the measured covariance matrix, the MENT algorithm converged to the linear-scale profiles in one iteration and the log-scale profiles in another two iterations, terminating at a mean absolute error of $10^{-4}$ per profile. Fig. 5 displays the measured and simulated beam profiles. We would like to highlight the agreement of the profiles down to the $10^{-3}$ level; with improved dynamic range in the wire scanner profiles, MENT may be able to study halo formation processes in two- or four-dimensional phase space. (We do not know if the measured asymmetric tails are real particles or if they are due to background or cross-talk between wires.)
Figure 5: Measured (black) vs. predicted (red) density on each wire. The measurement index is printed on the top left of each subplot.
Fig. 6 shows the marginal projections of the reconstructed four-dimensional phase space distribution.
Figure 6: Two-dimensional projections of the reconstructed four-dimensional phase space distribution. Contours vary linearly from $10^{-2}$ to $10^{0}$. A Gaussian filter with $\sigma=1$ pixel is applied to each image.
The nonstandard painting method used in this study ideally generates a uniform density beam with strong linear cross-plane correlations; space charge and other effects blur these correlations and generate a nonuniform density. The remaining linear correlations would be significant if this painting scheme were used for neutron production. In Fig. 7, we see that the average beam density is larger than expected from an uncorrelated beam with the same $x$-$x^{\prime}$ and $y$-$y^{\prime}$ distribution; the accelerator optics would need to be optimized to meet the target specifications. Further examination of the distribution is beyond the scope of this paper.
Figure 7: Ratio between the peak density of the reconstructed beam and a decorrelated version of the same beam ($f(x,x^{\prime},y,y^{\prime})\rightarrow f(x,x^{\prime})f(y,y^{\prime})$) as a function of position in the RTBT. Wire scanner locations are gray vertical lines.
The reconstruction quality depends on the accuracy of the forward model, including the measurement process. Our model represents the beamline as a $4\times 4$ transfer matrix:
where $\bm{\mathbf{M}}_{xx}$ transforms the $x$-$x^{\prime}$ coordinates, $\bm{\mathbf{M}}_{yy}$ transforms the $y$-$y^{\prime}$ coordinates, and $\bm{\mathbf{M}}_{xy}$ and $\bm{\mathbf{M}}_{yx}$ couple the horizontal and vertical planes. We assume $\bm{\mathbf{M}}_{xy}=\bm{\mathbf{M}}_{yx}=\bm{\mathbf{0}}$ because there are no coupled elements in the RTBT. The following are potential sources of error. (i) The true quadruple coefficients may not be equal to the readback values. We assume these values are accurate because measured twiss parameters typically agree with design values [13]; in future studies, orbit-difference measurements could use BPM data to confirm the linear model accuracy. (ii) Tilted quadrupoles would generate off-block-diagonal terms in the transfer matrix. We expect skew quadrupole components to be negligible, as they can be measured using BPM data [16] and would result in a tilted image on the target during neutron production. (iii) An incorrect beam energy would cause a systematic error in the quadrupole focusing strengths. We ignore this source of error because the beam energy is known to within 0.1% from time of flight measurements in the ring [17]. (iv) The quadrupole coefficients depend on the particle energy—each particle sees a slightly different transverse focusing force—but our model assumes the beam is monoenergetic. Strong transverse-longitudinal correlations in the bunch could amplify the error. The energy spread in the SNS is only a few MeV on top of a 1 GeV synchronous particle energy, so we expect that our model is sufficient. (v) The model does not include space charge. The tune shift over one turn in the ring is approximately 3%, too small to make a difference in the RTBT wire scanner region [13]. (vi) The wire scanner measurements assume the distribution has no pulse-to-pulse variation. Back-to-back measurements show almost no change in the profiles. An electron scanner [8] could also be used to validate this assumption in a future experiment.
It is important to note that tomography is an inverse problem with no unique solution, even with a perfect forward model and noiseless data. MENT returns the most reasonable inference from the data and prior but does not guarantee that the data are informative. (MENT returns the most reasonable solution even in the absence of data.) No existing high-dimensional reconstruction algorithms report uncertainty, so we rely heavily on reconstructions from fake data sets generated by simulated beams that reflect realistic conditions at the reconstruction point. To this end, we generated a test beam using a PyORBIT [18] simulation with parameters qualitatively similar to our experiment. (The simulation was reported in [2].) We rescaled the distribution to match our measured covariance matrix and transported it to the wire scanners using the same transfer matrices as in the experiment. Fig. 8 shows the test results. The reconstructed two-dimensional marginal distributions agree with the true two-dimensional distributions in both the linear correlations (as expected) and nonlinear correlations between planes. This test gives us confidence that MENT generates a reliable inference from the data but also highlights the need for robust uncertainty quantification techniques in real experiments.
Figure 8: Reconstruction from fake data. The initial distribution was generated from a PyORBIT simulation with a linear transformation applied to match the $x$-$x^{\prime}$ and $y$-$y^{\prime}$ covariance matrices to the measured values at the reconstruction point (see Fig. 4). The beam was transported to each wire scanner using the same transfer matrices as in the experiment. (a) Reconstructed distribution in normalized coordinates. Contours vary linearly from $10^{-2}$ to $10^{0}$. (b) Simulated (red) vs. “measured” (black) profiles.
IV Conclusion
We have demonstrated four-dimensional phase space tomography using one-dimensional measurements in a high-power hadron ring. We performed the reconstruction using the MENT algorithm with a Gaussian prior defined by the best-fit covariance matrix. The reconstructed distribution reproduces the measured profiles with the same dynamic range as the measurement devices. Test reconstructions with a known ground truth indicate a reasonably small reconstruction uncertainty.
An immediate use for SNS operations is to predict the two-dimensional beam density on the mercury spallation target [19, 20] or planned Second Target Station [21]. A straightforward extension in the SNS is to measure the four-dimensional density as a function of time by extracting the beam on different turns during injection. Such measurements could serve as benchmarks for computer simulations of intense, coupled beam dynamics in hadron rings. Finally, an electron scanner [8] could provide turn-by-turn profiles in the ring; it is unclear if one could use this data to infer the two- or four-dimensional phase space density.
V Acknowledgments
Appendix A Reconstruction method
@@
We aim to infer the probability density function $f(\bm{\mathbf{v}})$ from a set of one-dimensional projections, where $\bm{\mathbf{v}}=(v_{1},\dots,v_{N})$ is the phase space coordinate vector. (In four-dimensional phase space, $\bm{\mathbf{v}}=(x,x^{\prime},y,y^{\prime})$, where $x$ and $y$ are the positions and $x^{\prime}$ and $y^{\prime}$ are the momenta.) We assume the $i$th projection is measured after a linear transformation
@@
where $\bm{\mathbf{M}}_{i}$ is an $N\times N$ matrix representing the accelerator transport between the reconstruction point and the measurement device. We measure the one-dimensional projected density $g_{i}(w_{ij})$, where $w_{ij}$ is the $j$th dimension of $\bm{\mathbf{w}}_{i}$. Thus, the distribution must satisfy the following constraints:
@@
Entropy maximization (MaxEnt) provides a logically consistent way to incorporate prior information in the reconstruction and identify a unique solution from the many that satisfy Eq. (4). MaxEnt updates a prior distribution $f_{*}(\bm{\mathbf{v}})$ to a posterior by maximizing the relative entropy
@@
where $\lambda_{i}(w_{ij})$ are unknown Lagrange multiplier functions. The MENT algorithm uses a nonlinear Gauss-Seidel relaxation method to find $\lambda_{i}$ such that the distribution in Eq. (6) satisfies the constraints in Eq. (4). We use a sampling-based method to implement the Gauss-Seidel iterations; this will be reported in a separate paper.
An advantage of the MENT approach is its incorporation of external information through $f_{*}(\bm{\mathbf{v}})$. We note that the $N\times N$ covariance matrix ${\bm{\mathbf{\Sigma}}}=\langle\bm{\mathbf{v}}\bm{\mathbf{v}}^{T}\rangle$ is overdetermined by the measured beam profiles. Thus, we assume a Gaussian prior of the form
@@
where $\hat{\bm{\mathbf{\Sigma}}}$ is the best-fit covariance matrix determined by the following procedure. Under the symplectic linear transformation $\bm{\mathbf{M}}_{i}$, $\bm{\mathbf{\Sigma}}$ transforms as
@@
Eq. (8) linearly relates the measured rms beam size $\eta_{i}=\langle w_{ij}^{2}\rangle$ to the lower-triangular elements $\bm{\mathbf{\sigma}}$ of the initial covariance matrix. Stacking the measurements into a single vector $\bm{\mathbf{\eta}}=(\eta_{1},\eta_{2},\dots)$ gives the following system of equations:
$\bm{\mathbf{A}}\bm{\mathbf{\sigma}}=\bm{\mathbf{\eta}}.$ (9)
We generate a solution $\hat{\bm{\mathbf{\sigma}}}$ (and therefore $\hat{\bm{\mathbf{\Sigma}}}$) to Eq. (9) using ordinary linear least squares (LLSQ) [22]:
@@
With the LLSQ fit in hand, we may run the MENT algorithm to reconstruct the beam distribution in normalized phase space coordinates,

### v2

# Four-dimensional phase space tomography from one-dimensional measurements of a hadron beam

@@ Abstract
In this paper, we use one-dimensional measurements to infer the four-dimensional phase space density of an accumulated proton beam in the Spallation Neutron Source (SNS) accelerator. The reconstruction was performed by maximizing the distribution’s entropy subject to the measurement constraints, and thus represents the most conservative inference from the data. The reconstructed distribution reproduces the measured profiles down to the noise level, and simulations indicate that the problem is reasonably well-constrained. Similar measurements could serve as benchmarks for beam dynamics simulations in the SNS or hadron accelerators.
@@ I Introduction
High-dimensional phase space measurements provide valuable information about the beam dynamics in an accelerator; however, conventional diagnostics can only measure the distribution’s low-dimensional projections. Reconstructing the phase space distribution from these projections is referred to as phase space tomography.
Let us focus on the specific case of four-dimensional phase space. Recent experiments have demonstrated four-dimensional tomography from two-dimensional measurements in electron accelerators [1, 2, 3], as well as in an H- linac using laser wires [4]. However, two-dimensional measurements are not always available. After the initial acceleration stages in high-power accelerators, the primary diagnostics are wire scanners, which record the beam density on a one-dimensional axis. It is not a priori obvious whether a realistic number of one-dimensional views can tightly constrain a four-dimensional distribution, nor whether the required views can be obtained in an accelerator.
To our knowledge, the only published 4D-from-1D reconstruction was by Minerbo, Sander, and Jameson in 1981 [5]. The authors fit a four-dimensional distribution to 10 profiles. The sole figure in the paper compares the reconstructed $x$-$x^{\prime}$ and $y$-$y^{\prime}$ distributions to direct measurements. This comparison says nothing about the accuracy in the four-dimensional space; further study is needed to determine the usefulness of four-dimensional tomography when only one-dimensional measurements are available.
To this end, in this paper, we use one-dimensional measurements to infer the four-dimensional phase space distribution of an accumulated proton beam in the Spallation Neutron Source (SNS) accelerator. We measure the beam using four wire scanners in a transfer line between the proton accumulator ring and the spallation target. Following [5], we reconstruct the distribution using the method of maximum relative entropy. The reconstructed distribution reproduces the measured profiles down to the wire scanner noise level.
As a form of uncertainty quantification, we simulate the reconstruction with fake data generated by various ground truth distributions. We find that the two-dimensional marginal distributions ($x$-$x^{\prime}$, $y$-$y^{\prime}$, $x$-$y$, $x$-$y^{\prime}$, $y$-$x^{\prime}$, $x^{\prime}$-$y^{\prime}$) are reasonably well constrained by the data. We also find that these low-dimensional views can obscure certain higher-dimensional features, such as four-dimensional hollowing. We conclude that our set of optics is sufficient for typical beams in the SNS, but that different optics or additional measurements may be needed to reconstruct arbitrary distributions in the full four-dimensional phase space.
Organization: In Section II, we review the method of maximum relative entropy. In Section III, we describe the SNS accelerator, diagnostics, and experimental details. In Section IV, we report the reconstruction results and perform a suite of reconstructions from fake data. Section VI concludes.
II Reconstruction algorithm
@@
We use the method of maximum relative entropy to infer the phase space distribution from its one-dimensional projections. Denote the distribution function as $f(\bm{\mathbf{v}})$, where $\bm{\mathbf{v}}=(v_{1},\dots,v_{N})$ is the phase space coordinate vector. (In four-dimensional phase space, $\bm{\mathbf{v}}=(x,x^{\prime},y,y^{\prime})$, where $x$ and $y$ are the positions and $x^{\prime}$ and $y^{\prime}$ are the momenta.) We assume the $i$th projection is measured after a linear transformation
@@
where $\bm{\mathbf{M}}_{i}$ is an $N\times N$ matrix representing the accelerator transport between the reconstruction point and the measurement device, and $\mathbf{w}_{i}=(w_{i1},\dots,w_{iN})$ is the transformed coordinate vector. We measure the projected density $g_{i}(w_{ij})$. Thus, the distribution must satisfy the following constraints:
@@
To select a distribution compatible with Eq. (2), we update a prior distribution $f_{*}(\bm{\mathbf{v}})$ to a posterior by maximizing the relative entropy
@@
where $\lambda_{i}(w_{ij})$ are Lagrange multiplier functions. The MENT algorithm uses a nonlinear Gauss-Seidel relaxation method to find $\lambda_{i}$ such that the distribution in Eq. (4) satisfies the constraints in Eq. (2). We use the particle-based version of MENT described in [6].
An advantage of the maximum entropy method is the ability to incorporate external information through the prior $f_{*}(\bm{\mathbf{v}})$. Following [5], because the covariance matrix ${\bm{\mathbf{\Sigma}}}=\langle\bm{\mathbf{v}}\bm{\mathbf{v}}^{T}\rangle$ is overdetermined by the measured second-order moments, we consider $\bm{\mathbf{\Sigma}}$ to be prior information. For best-fit covariance matrix $\hat{\bm{\mathbf{\Sigma}}}$, we define the following Gaussian prior:
@@
(The Gaussian distribution maximizes the absolute entropy subject to a known covariance matrix.) There is no requirement to define the prior in this way, but it is a reasonable choice in accelerator physics when the covariance matrix is known.
We use the following standard recipe to fit the covariance matrix to the measured second-order moments. Under the symplectic linear transformation $\bm{\mathbf{M}}_{i}$, $\bm{\mathbf{\Sigma}}$ transforms as
@@
Eq. (6) linearly relates the measured second-order moments $\eta_{i}=\langle w_{ij}^{2}\rangle$ to the lower-triangular elements $\bm{\mathbf{\sigma}}$ of the initial covariance matrix. Stacking the measurements into a single vector $\bm{\mathbf{\eta}}=(\eta_{1},\eta_{2},\dots)$ gives the following system of equations:
$\bm{\mathbf{A}}\bm{\mathbf{\sigma}}=\bm{\mathbf{\eta}},$ (7)
where $\bm{\mathbf{A}}$ is determined by the transfer matrix elements. We generate a solution $\hat{\bm{\mathbf{\sigma}}}$ (and therefore $\hat{\bm{\mathbf{\Sigma}}}$) to Eq. (7) using ordinary linear least squares (LLSQ) [7]:
@@
After computing the LLSQ solution and resulting prior (Eq. (5)), we run the MENT algorithm to reconstruct the beam distribution in normalized coordinates,